# Guangyu Zhu, Yiqin Deng, , Xianhao Chen, , Haixia Zhang, , Yuguang Fang, , Tan F. Wong

### v1

@@ I Introduction
Federated learning (FL), a privacy-preserving distributed ML technique enables EDs to collaboratively learn a global ML model without sharing their raw data, consequently reducing the requirement of communication bandwidth between EDs and a central server as well. The intuitive idea of privacy-preserving distributed ML was investigated independently by some researchers Xu et al. [5, 6, 7], Shokri et al. [8], Gong et al. [9]. McMahan et al. [10] first constructed a distributed ML framework based on decentralized datasets held by different users, and claimed their algorithm as FL. The original FL algorithm updates ML models on EDs locally, and aggregates the updated ML models to obtain the global model at the central server remotely, in contrast to the traditional ML requiring EDs to upload their private data to where the ML model is trained. Since FL retains users’ private data on the EDs without sharing raw data with the server, the server only aggregates users’ locally updated models, as the privacy of end users is naturally preserved to some extent. Besides, as all training processes are performed by the EDs, the computation and communication capacities of EDs can also be utilized. FL provides various advantages, such as data privacy preservation, reduced communication latency, and enhanced learning performance [2]. However, pushing all training workload to EDs sometimes is impracticable. Training complex ML models often consumes unacceptable amount of training memory and computing power on Internet of Things (IoT) devices with limited communication and computing resources, and incurs intolerable latency. Besides, the ML model size sometimes reaches GBs (e.g., GPT-1 has 0.12 billion parameters, GPT-2 has 1.5 billion parameters, whereas GPT-3 has more than 175 billion parameters), resulting in a significant communications burden duo to the need of frequently uploading local ML models.
To cope with the dilemma between insufficient ED resources and complicated ML models, we leverage another ML technique, called split federated learning (SFL) [11], which introduces model splitting from split learning (SL) [12] to FL. The split federated learning framework was proposed by Thapa et al. [11] in 2020 to integrate federated learning with split learning, shown in Figure 1. Under SFL framework, the main server helps the training process for each client, and the Fed server is tasked with the aggregation of all locally updated models. The SFL approach presents a compelling advantage for resource-constrained environments, since the main server undertake parts of local training workload.
@@
From the conventional FL and SFL implementations across heterogeneous devices, we have identified several limitations that can significantly impede the system’s overall performance and efficiency. In synchronous FL and SFL, the aggregation process is inherently constrained by the pace of the slowest participant due to the necessity for the server to collect updated local gradients from all selected EDs. This synchronicity results in a scenario where devices with abundant computational resources are rendered idle as they wait for the transmission of local models from less capable devices, often referred to as “straggler”. This inefficiency is primarily attributed to the varying and unpredictable communication and computational capabilities of the heterogeneous EDs. Therefore, addressing resource heterogeneity (RH), also called the system heterogeneity problem in FL, is critical to enhancing the efficiency and overall feasibility of SFL.
In this paper, we propose a novel efficient split federated learning algorithm (ESFL) to boost training efficiency and performance considering the privacy-preserving constraint. The cornerstone of ESFL lies in its strategic utilization of heterogeneity in system resources and device capabilities. Unlike standard synchronous FL and SFL, our approach involves the server in the training phase but dynamically adjusts the distribution of user-side workload and server-side resources. The ESFL algorithm thereby capitalizes on the intrinsic resource variability across EDs to optimize ML outcomes and system efficiency.
@@ II Related Works
There exist quite extensive research on splitting learning and federated learning in the current literature. In this section, we will concentrate on closely related works to review. In [13], Ang et al. offered a robust architecture for FL to increase communication efficiency by reducing transmission noise in wireless networks. To address the communications overhead in FL, Wang et al. [14] developed an algorithm, named Communication Mitigated Federated Learning (CMFL), to eliminate irrelevant local model updates that were trained over users’ biased data. It should be noted that the aforementioned two methods only address communication efficiency in FL, as all training processes are executed by the user-side EDs. FedMMD was proposed in [15] to reduce communication and computing costs during the local training process in FL using a two-stream model with Maximum Mean Discrepancy (MMD) to replace the local training for a single model in FedAVG [10]. However, while this method reduces the number of communication rounds, it concurrently increases the user-side workload due to the need of computing the MMD loss. Wenqi et al. [16] address both communication and computing resource heterogeneities in wireless FL by providing a joint device scheduling and resource allocation strategy. Despite this integrated approach, the limitation of FL is still evident, as the user-side workload cannot be reduced.
@@
In [11], Thapa et al. designed a novel framework, called splitfed learning (SFL), to take advantage of parallel training among different users in FL and the model splitting in SL to reduce the computing workload on EDs. In [17], Gao et al. implemented SFL and evaluated the performance on IoT devices. In [18], Lin et al. developed an efficient parallel split learning algorithm by applying the last-layer gradient aggregation to reduce communication and computing overheads in SL. In [19], Wu et al. designed a resource allocation strategy for cluster-based SFL. In [20], Kim et al. devised a bargaining game to negotiate the cut layer in personalized parallel SL. All aforementioned literature on SFL overlook the crucial aspect of the joint consideration of both user-side resource and workload heterogeneities. This oversight is evidenced by the fact that the cut layer remains the same at all EDs. However, it is obvious that adjusting cut layers can significantly change user-side workload distribution.
@@
To reduce the computing and communication workload at EDs when considering RH in FL, Sattler et al.  [21] designed a novel model compression algorithm to extend the commonly used top-k gradient sparsification to FL to compress both model uploading and downloading. However, since model compression inevitably results in performance degradation, it is advisable to maintain the integrity of the model architecture throughout the training process. With the emergence of edge computing [22], utilizing the computing resources at the both the central server and edges can be leveraged to reduce workload at EDs. For instance, in [23], Wang et al. proposed to enable EDs to collaborate with edge nodes by exchanging model parameters to reduce the user-side workload, and to apply deep reinforcement learning to optimize the operations of mobile edge computing, caching, and communications. Several researchers attempted to leverage edge nodes to assist FL by fully uploading local training tasks to trusted edge nodes to reduce the user-side workload [24]. Unfortunately, the trustworthiness of training edge nodes is often hard to guarantee. Our ESFL algorithm introduce an integrated strategy for the allocation of server-side computing resource and user-side training workload. This approach is designed to accommodate the variations in computing and communication resources inherent in heterogeneous EDs. We will present the detailed design next.
@@ III-A Motivation
While SFL only address the resource constraints at EDs, it does not account for the variations in data distribution (DH) and resource availability (resource heterogeneity or RH) inherent in EDs. Within a SFL framework, clients or EDs are required to partition their models following an identical structure, leading to a scenario where the communication and computational workload on the client side is influenced solely by the volume of data. In contrast, our Efficient Split Federated Learning (ESFL) takes a holistic approach, considering both the client-side workload and the server-side computing resource allocation, which is designed to mitigate “stragglers” problem in FL, thereby enhancing the training efficiency of SFL.
@@ III-B ESFL framework
    Model splitting and resource allocation: The server first acquires the information of the selected EDs, including data amount, and available communication and computing resources. Then, the server jointly splits the ML model into user-side models and server-side models (the so-called cut layer decision), and allocates adequate server-side resources based on the users’ information and the ML model structure for the selected user.
@@
¡¡¡¡¡¡¡ HEAD
@@ III-B1 Split Training
The pseudocode for our ESFL is shown in Algorithm 1 and Algorithm 2. The split training is shown in Algorithm 1, while Step $2$ is to let users send activations of the cut layer together with the label to the server. ======= The pseudocode for ESFL is shown in Algorithm 1 and Algorithm 2. The Split training is shown in Algorithm 1, where Step $2$ is to let users send activations of the cut layer together with the label to the server. The detail implementing steps of ESFL are shown in Algorithm 2. ¿¿¿¿¿¿¿ 8fed39b787cde27c064d35637df79221d613d121
@@ III-B2 Federated Aggregation
¡¡¡¡¡¡¡ HEAD In ESFL, we leverage FedAVG algorithm [10] to aggregate multiple user-side and server-side updated ML models and add a scaling factor $\eta$. We call this aggregation method as federated aggregation, which mitigates the training oscillation by leveraging the global ML model in the previous round, compared with FedAVG, which only aggregates current-round updated local ML model to generate next-round global ML models. The architecture of federated aggregation is shown in Figure 3. We define the whole ML model trained by user $i$ in round $r$ as $\mathcal{w}_{i}^{r}$ and the global model in round $r$ is $\mathcal{\textbf{W}}^{r}$. The update of the global ML model in round ${r+1}$ is ======= In ESFL, we leverage the FedAVG algorithm [10] to aggregate multiple user-side and server-side updated ML models using $\eta$ as a step size factor. We call this aggregation method federated aggregation, which mitigates training oscillations by leveraging the global ML model in the previous round. This differ from FedAVG which only aggregates the current-round updated local ML models to generate the next-round global ML model. The architecture of federated aggregation is shown in Figure 3. We define the whole ML model trained by user $i$ in round $r$ as $\mathcal{w}_{i}^{r}$ and the global model in round $r$ is $\mathcal{\textbf{W}}^{r}$. The update of the global ML model in round ${r+1}$ is ¿¿¿¿¿¿¿ 8fed39b787cde27c064d35637df79221d613d121
@@
The whole training model is split into two parts, namely, user-side model $\textbf{W}_{u,i}^{r}$ trained by a user-side ED and the server-side model $\textbf{W}_{s,i}^{r}$ trained by a virtual server $v_{i}$. All virtual servers are either virtual machines or containers at the central server, which are allocated with corresponding computing and communication resources according to the model training demands at this round. Thus, after concatenating the user-side ML model and the server-side ML model, for each selected user, the concatenated ML model will have the same architecture as the global ML model. The updated concatenated ML model for user $i$ in round $r+1$ is
@@
1 Initialize the global model parameters W;
@@ 3      The server randomly selects users $S$ to participate the $r$-th round training;
5      The server allocates its resources $C^{*}$ and $B^{*}$, and determine cut layer $l^{*}$ based on $c^{*}$ and $b^{*}$;
@@ III-B3 Communication Model
where the communication bandwidth allocated to the user $i$ is $B_{i}$, and the total available bandwidth is $B$ where $\sum_{i=1}^{S}B_{i}\leq B$, $\gamma_{i}^{u}$ is the uplink channel gain and $\gamma_{i}^{d}$ is the downlink channel gain for user $i$, $P^{u}_{i}$ and $P^{d}_{i}$ are the uplink and downlink transmission powers, respectively, when the resource block $B_{i}$ is used, where $P^{u}_{i}$, $P^{d}_{i}$, $\gamma_{i}^{u}$, and $\gamma_{i}^{d}$ are predetermined constant values. $N_{0}$ is the noise power density. We assume that the transmission environment is stationary during one training round. For example, EDs can be cameras in smart homes, whose deployment locations remain fixed for a relatively long period.
@@ III-B4 Workload and Resource Allocation
To reduce the training workload on user-side EDs, we split ML training into local training and server ML training. Thus, the next problem is how to appropriately determine user-side communication and computing workload, alongside the strategic allocation of computing resources on the server to different virtual servers. Due to the varying sizes of data collected by different users/EDs and the heterogeneous available resources on EDs, simply assigning the same amount of resources and randomly choosing the cut layers for different users will not help the training time and the performance. To maximize the training efficiency, one should address a joint optimization of workload and resource.
@@
where $b_{i}^{U}$ and $b_{i}^{D}$ are the uplink and downlink data rate for user $i$ as defined in (5). We denote the index of the cut layer for user $i$ as $l_{i}$ and $\epsilon_{i}$ as the number of local epochs for user $i$. The uploading and downloading workload are assumed to be $M^{l}_{i}$, equal to the user-side model size of user $i$. Since each selected user applies the same local dataset to train ML model and the idle resource is also stationary, to simplify the following optimization problem, in this study, we assume that the training times of all epochs are constant in each round of training time.
Since both user-side EDs and the server participate in each epoch model updating, one-epoch training time contains four parts, namely, local/user-side computing time $t_{i}^{c}$, uploading time for activation $t_{i}^{b}$, remote/server-side computing time $t_{i}^{C}$, and downloading time for updated activation $t_{i}^{B}$. Thus, one epoch time can be denoted as
$\begin{gathered}T^{e}_{i}=t^{c}_{i}+t^{b}_{i}+t^{C}_{i}+t^{B}_{i}\\ =\frac{w^{c}_{i}}{c_{i}}+\frac{w^{b}_{i}}{b_{i}}+\frac{W^{c}_{i}}{C_{i}}+\frac{w^{b}_{i}}{B_{i}}\\ =\frac{\sum_{l}{D^{l}_{c}x^{l}_{i}\cdot n_{i}}}{c_{i}}+\frac{\sum_{l}{D^{l}_{b}x^{l}_{i}\cdot n_{i}}}{b^{u}_{i}}+\\ \frac{(D-\sum_{l}{D^{l}_{c}x^{l}_{i}})\cdot n_{i}}{C_{i}}+\frac{\sum_{l}{D^{l}_{b}x^{l}_{i}}\cdot n_{i}}{b^{d}_{i}}.\\ \end{gathered}$ (8)
The cut layer for user $i$ is $l_{i}\in[1,2,...,L]$, where the whole ML neural network is composed of $L$ layers, and the server-side model and the user-side model for user $i$ are split at the $l_{i}$-th layer. To simplify the notation of workloads, we denote $x^{l}_{i}$ as a $[0,1]$ vector, where $\sum_{l}x^{l}_{i}=1$. The total computing workload for training one sample is $D$ and the user-side computing workload for user $i$ is $w^{c}_{i}=\sum_{l}{D^{l}_{c}x^{l}_{i}\cdot n_{i}}$, where $D^{l}$ is the computing workload for one training sample for user-side EDs when the cutting layer is $l$. The upload and download data size for user $i$ is $w_{i}^{b}=\sum_{l}{{D^{l}_{b}x^{l}_{i}}\cdot n_{i}}$, where $D^{l}_{b}$ is the size of activation data for one training sample while the cut layer is $l$. For user $i$, computing capability is $c_{i}$, uplink transmission rate is $b^{u}_{i}$, and downlink transmission rate is $b^{d}_{i}$. The computing resource that the server allocates to user $i$ is $C_{i}$.

### v2

@@ I Introduction
Federated learning (FL), a privacy-preserving distributed ML technique enables EDs to collaboratively learn a global ML model without sharing their raw data, consequently reducing the requirement of communication bandwidth between EDs and a central server as well. The intuitive idea of privacy-preserving distributed ML was investigated independently by some researchers Xu et al. [5, 6, 7], Shokri et al. [8], Gong et al. [9]. McMahan et al. [10] first constructed a distributed ML framework based on decentralized datasets held by different users, and coined their algorithm as FL. The original FL algorithm updates ML models on EDs locally, and aggregates the updated ML models to obtain the global model at a central server remotely, in contrast to the traditional ML requiring EDs to upload their private data to where the ML model is trained. Since FL retains users’ private data on EDs without sharing raw data with the server, the server only aggregates users’ locally updated models, and thus the privacy of end users is naturally preserved to some extent. Besides, as all training processes are performed by the EDs, the computation and communication capacities of EDs can also be utilized. FL provides various advantages, such as data privacy preservation, reduced communication latency, and enhanced learning performance [2]. However, pushing all training workload to EDs sometimes is impracticable. Training complex ML models often consumes unacceptable amount of training memory and computing power on Internet of Things (IoT) devices with limited communication and computing resources, and incurs intolerable latency. Besides, the ML model size sometimes reaches GBs and even TBs(e.g., GPT-1 has 0.12 billion parameters, GPT-2 has 1.5 billion parameters, whereas GPT-3 has more than 175 billion parameters), resulting in a significant communications burden duo to the need of frequently uploading local ML models.
To cope with the dilemma between insufficient ED resources and complicated ML models, we leverage another ML technique, called split federated learning (SFL) [11], which introduces model splitting from split learning (SL) [12] to FL. The SFL framework was proposed by Thapa et al. [11] in 2020 to integrate federated learning with split learning, shown in Figure 1. Under SFL framework, the main server helps the training process for each client, and the Fed server is tasked with the aggregation of all locally updated models. The SFL approach presents a compelling advantage for resource-constrained environments, since the main server undertake parts of local training workload.
@@
From the conventional FL and SFL implementations across heterogeneous devices, we have identified several limitations that can significantly impede the system’s overall performance and efficiency. In synchronous FL and SFL, the aggregation process is inherently constrained by the pace of the slowest participant due to the necessity for the server to collect updated local gradients from all selected EDs. This synchronicity results in a scenario where devices with abundant computational resources are rendered idle as they wait for the transmission of local models from less capable devices, often referred to as “stragglers”. This inefficiency is primarily attributed to the varying and unpredictable communication and computational capabilities of heterogeneous EDs. Therefore, addressing resource heterogeneity (RH), also called the system heterogeneity in FL, is critical to enhancing the efficiency and overall feasibility of SFL.
In this paper, we propose a novel efficient split federated learning algorithm (ESFL) to boost training efficiency and performance by considering the privacy-preserving constraints. The cornerstone of ESFL lies in its proactive strategic utilization of heterogeneity in system resources and device capabilities. Unlike standard synchronous FL and SFL, our approach involves the server in the training phase but dynamically adjusts the distribution of user-side workload and server-side resources. The ESFL algorithm thereby capitalizes on the intrinsic resource variability across EDs to optimize ML outcomes and system efficiency.
@@ II Related Works
There exist quite extensive research on splitting learning (SL) and federated learning (FL) in the current literature. In this section, we will concentrate on closely related works to review. In [13], Ang et al. offered a robust architecture for FL to increase communication efficiency by reducing transmission noise in wireless networks. To address the communications overhead in FL, Wang et al. [14] developed an algorithm, named Communication Mitigated Federated Learning (CMFL), to eliminate irrelevant local model updates that were trained over users’ biased data. It should be noted that the aforementioned two methods only address communication efficiency in FL, as all training processes are executed by users at EDs. FedMMD was proposed in [15] to reduce communication and computing costs during the local training process in FL using a two-stream model with Maximum Mean Discrepancy (MMD) to replace the local training for a single model in FedAVG [10]. However, while this method reduces the number of communication rounds, it concurrently increases the user-side workload due to the need of computing the MMD loss functions. Shi et al. [16] address both communication and computing resource heterogeneities in wireless FL by providing a joint device scheduling and resource allocation strategy. Despite this integrated approach, the limitation of FL is still evident, as the user-side workload cannot be reduced. There exist also some other innovative approaches to ease the work load for communications and computing for federated learning. Watanabe et al. [17] and Chen et al. [18] leveraged wireless mesh networks to either facilitate communications or reduce communications traffic. Guo et al. [19, 20] utilized edge nodes and federated reinforcement learning to help resource-constrained D2D devices in industrial IoT systems and 5G networks, which can address the device heterogeneity issue to some extent.
@@
In [11], Thapa et al. designed a novel framework, called split federated learning (SFL), to take advantage of parallel training among different users in FL and the model splitting in SL to reduce the computing workload on EDs. In [21], Gao et al. implemented SFL and evaluated the performance on IoT devices. In [22], Lin et al. developed an efficient parallel split learning algorithm by applying the last-layer gradient aggregation to reduce communication and computing overheads in SL. In [23], Wu et al. designed a resource allocation strategy for cluster-based SFL. In [24], Kim et al. devised a bargaining game to negotiate the cut layer in personalized parallel SL. All aforementioned literature on SFL overlook the crucial aspect of the joint consideration of both user-side resource and workload heterogeneities. This oversight is evidenced by the fact that the cut layer remains the same at all EDs. However, it is obvious that adjusting cut layers can significantly change user-side workload distribution.
@@
To reduce the computing and communication workload at EDs when considering RH in FL, Sattler et al.  [25] designed a novel model compression algorithm to extend the commonly used top-k gradient sparsification to FL to compress both model uploading and downloading. However, since model compression inevitably results in performance degradation, it is advisable to maintain the integrity of the model architecture throughout the training process. With the emergence of edge computing [26], utilizing the computing resources at the both the central server and edges can be leveraged to reduce workload at EDs. For instance, in [27], Wang et al. proposed to enable EDs to collaborate with edge nodes by exchanging model parameters to reduce the user-side workload, and to apply deep reinforcement learning to optimize the operations of multi-access edge computing (MEC), caching, and communications. Several researchers attempted to leverage edge nodes to assist FL by fully uploading local training tasks to trusted edge nodes to reduce the user-side workloads [28]. Unfortunately, the trustworthiness of training edge nodes is often hard to guarantee. Our ESFL algorithm introduce an integrated strategy for the allocation of server-side computing resource and user-side training workloads. This approach is designed to accommodate the variations in computing and communication resources inherent in heterogeneous EDs. We will present the detailed design next.
@@ III-A Motivation
While SFL only address the resource constraints at EDs, it does not account for the variations in data distribution (DH) and resource availability (resource heterogeneity or RH) inherent in EDs. Within the SFL framework, clients or EDs are required to partition their models following an identical structure, leading to a scenario where the communication and computational workload on the client side is influenced solely by the volume of data. In contrast, our Efficient Split Federated Learning (ESFL) takes a holistic approach, considering both the client-side workload and the server-side computing resource allocation, which is designed to mitigate “stragglers” problem in FL, thereby enhancing the training efficiency of SFL. It is worth noting that our method is a scheme for joint optimization of resource and workload allocation, which can be integrated with any user selection algorithms presently existed in FL. The ESFL framework demonstrates the capability to increase training efficiency across the board, independent of the particular user selection algorithm integrated in the framework. This underscores the inherent adaptability and effectiveness of our proposed scheme in enhancing the training efficacy of FL.
@@ III-B ESFL framework
    Model splitting and resource allocation: The server first acquires the information of the selected EDs, including data amount, available communication, and computing resources. Then, the server jointly splits the ML model into user-side models and server-side models (the so-called cut layer decision) and allocates adequate server-side resources based on the users’ information and the ML model structure for the selected user.
@@
TABLE I: Summary of Important Notations of ESFL
66 Notation Description $S$ The set of selected users $R$ The total training rounds $\epsilon_{i}$ The number of local training epochs of user $i$ $\textbf{W}_{u,i}^{r,e}$ The user-side model for user $i$ at epoch $e$ at round $r$ $\textbf{W}_{s,i}^{r,e}$ The server-side model for user $i$ at epoch $e$ at round $r$ $\textbf{A}_{l_{i}}^{r,e}$ The activation calculated by the user $i$ at epoch $e$ at round $r$ using forward propagation algorithm $\eta$ The step size factor of the federated aggregation $N$ The number of total one-round training data samples $n_{i}$ The number of one-round training data samples for user $i$ $\textbf{W}^{r}$ The global model at round $r$ $\textbf{W}_{i}^{r}$ The cascaded local model generated by user $\textbf{W}_{u,i}$ and server-side model $\textbf{W}_{s,i}$ at round $r$ $b_{i}^{u}$ Uploading data rate of user $i$ $b_{i}^{d}$ Downloading data rate of user $i$ $B_{i}$ The bandwidth allocated to user $i$ $P^{u}_{i}$ The uplink transmission power of user $i$ $P^{d}_{i}$ The downlink transmission power of user $i$ $\gamma^{u}_{i}$ The uplink channel gain of user $i$ $\gamma^{d}_{i}$ The downlink channel gain of user $i$ $N_{0}$ The noise power density $T$ The total training time $T_{r,i}$ The $r$-th round training time for user $i$ $T_{i}^{U}$ The $r$-th round local model uploading time for user $i$ $T_{i}^{D}$ The $r$-th round global model downloading time for user $i$ $T_{i}^{e}$ The split training time at epoch $e$ for user $i$ $T^{agg}$ The federated aggregation time for the server at round $r$ $M^{l_{i}}_{i}$ The local model uploading/ downloading time for user $i$ $t_{i}^{e,c}$ The user-side computing time for user $i$ at epoch $e$ $t_{i}^{e,b}$ The uploading time of the user-side activation for user $i$ at epoch $e$ $t_{i}^{e,C}$ The server-side computing time for user $i$ at epoch $e$ $t_{i}^{e,B}$ The downloading time of the server-side updated activation for user $i$ at epoch $e$ $D_{c}^{l}$ The user-side computing workload of the cut layer $l$ for training one sample $D_{b}^{l}$ The user-side communication workload of the cut layer $l$ for training one sample $D$ The computing workload for training one data sample $c_{i}$ The local available computing resource for user $i$ $C_{i}$ The allocated server-side computing resource to user $i$ $x_{i}$ The cut layer indicator vector of user $i$ $s_{i}$ The available storage space for user $i$ $m_{i}$ The available memory space for user $i$
In what follows, we will provide more details for the whole procedure below.
@@ III-B1 Split Training
The pseudocode for our ESFL is shown in Algorithm 1 and Algorithm 2. The split training is shown in Algorithm 1, while Step $2$ is to let users send activations of the cut layer together with the label to the server.
@@ III-B2 Federated Aggregation
In ESFL, we leverage the FedAVG algorithm [10] to aggregate multiple user-side and server-side updated ML models using $\eta$ as a step size factor. We call this aggregation method federated aggregation, which mitigates training oscillations by leveraging the global ML model in the previous round. This differs from FedAVG which only aggregates the current-round updated local ML models to generate the next-round global ML model. The architecture of federated aggregation is shown in Figure 3. We define the whole ML model trained by user $i$ in round $r$ as $\mathcal{w}_{i}^{r}$ and the global model at round $r$ is $\mathcal{\textbf{W}}^{r}$. The update of the global ML model at round ${r+1}$ is
@@
The whole training model is split into two parts, namely, user-side model $\textbf{W}_{u,i}^{r}$ trained by a user-side ED and the server-side model $\textbf{W}_{s,i}^{r}$ trained by a virtual server $v_{i}$. All virtual servers are either virtual machines or containers at the central server, which are allocated with corresponding computing and communication resources according to the model training demands at this round. Thus, after concatenating the user-side ML model and the server-side ML model, for each selected user, the concatenated ML model will have the same architecture as the global ML model. The updated concatenated ML model for user $i$ at round $r+1$ is
@@
1
2Initialize the global model parameters W;
@@ 4      The server randomly selects users $S$ to participate the $r$-th round training;
6      The server allocates its resources $C^{*}$ and $B^{*}$ and determines cut layer $l^{*}$ based on $c^{*}$ and $b^{*}$;
@@ III-B3 Communication Model
where the communication bandwidth allocated to the user $i$ is $B_{i}$, and the total available bandwidth is $B$ where $\sum_{i=1}^{S}B_{i}\leq B$, $\gamma_{i}^{u}$ is the uplink channel gain and $\gamma_{i}^{d}$ is the downlink channel gain for user $i$, $P^{u}_{i}$ and $P^{d}_{i}$ are the uplink and downlink transmission powers, respectively, when the resource block $B_{i}$ is used, where $P^{u}_{i}$, $P^{d}_{i}$, $\gamma_{i}^{u}$, and $\gamma_{i}^{d}$ are predetermined constant values, and $N_{0}$ is the noise power density. We assume that the transmission environment is stationary during one training round. For example, EDs can be cameras in smart homes, whose deployment locations remain fixed for a relatively long period.
@@ III-B4 Workload and Resource Allocation
To reduce the training workload on user-side EDs, we split ML training into local training and server ML training. Thus, the next problem is how to appropriately determine user-side communication and computing workload, alongside the strategic allocation of computing resources on the server to different virtual servers. Due to the varying sizes of data collected by different users/EDs and the heterogeneous available resources on EDs, simply assigning the same amount of resources and randomly choosing the cut layers for different users will not help the training time and the training performance. To maximize the training efficiency, one should address a joint optimization of workload and resource.
@@
where $b_{i}^{U}$ and $b_{i}^{D}$ are the uplink and downlink data rate for user $i$ as defined in (5). We denote the index of the cut layer for user $i$ as $l_{i}$ and $\epsilon_{i}$ as the number of local epochs for user $i$. The uploading and downloading workload are assumed to be $M^{l}_{i}$, equal to the user-side model size of user $i$. Since each selected user applies the same local dataset to train ML model and the idle resource is also stationary, to simplify the following optimization problem, in this paper, we assume that the training times of all epochs are constant in each round of training time.
Since both user-side EDs and the server participate in each epoch model updating, one-epoch training time contains four parts, namely, local/user-side computing time $t^{e,c}$, uploading time for activation $t^{e,b}$, remote/server-side computing time $t^{e,C}$, and downloading time for updated activation $t^{e,B}$. Thus, one epoch time can be denoted as
$\begin{gathered}T^{e}_{i}=t^{e,c}_{i}+t^{e,b}_{i}+t^{e,C}_{i}+t^{e,B}_{i}\\ =\frac{w^{e,c}_{i}}{c_{i}}+\frac{w^{e,b}_{i}}{b_{i}}+\frac{W^{e,C}_{i}}{C_{i}}+\frac{w^{e,B}_{i}}{B_{i}}\\ =\frac{\sum_{l}{D^{l}_{c}x^{l}_{i}\cdot n_{i}}}{c_{i}}+\frac{\sum_{l}{D^{l}_{b}x^{l}_{i}\cdot n_{i}}}{b^{u}_{i}}+\\ \frac{(D-\sum_{l}{D^{l}_{c}x^{l}_{i}})\cdot n_{i}}{C_{i}}+\frac{\sum_{l}{D^{l}_{b}x^{l}_{i}}\cdot n_{i}}{b^{d}_{i}}.\\ \end{gathered}$ (8)
The cut layer for user $i$ is $l_{i}\in[1,2,...,L]$, where the whole ML neural network is composed of $L$ layers, and the server-side model and the user-side model for user $i$ are split at the $l_{i}$-th layer. To simplify the notation of workloads, we denote $x^{l}_{i}$ as an indicator function, where $\sum_{l}x^{l}_{i}=1$, and $x_{i}^{l}=1$ indicates that $l$ layer is chosen as the cut layer for user $i$ while $x_{i}^{l}=0$ indicates that $l$ layer is not selected as the cut layer. The total computing workload for training one sample is $D$ and the user-side computing workload for user $i$ is $w^{c}_{i}=\sum_{l}{D^{l}_{c}x^{l}_{i}\cdot n_{i}}$, where $D^{l}$ is the computing workload for one training sample for user-side EDs when the cut layer is $l$. The upload and download data size for user $i$ is $w_{i}^{b}=\sum_{l}{{D^{l}_{b}x^{l}_{i}}\cdot n_{i}}$, where $D^{l}_{b}$ is the size of activation data for one training sample while the cut layer is $l$. For user $i$, computing capability is $c_{i}$, uplink transmission rate is $b^{u}_{i}$, and downlink transmission rate is $b^{d}_{i}$. The computing resource that the server allocates to user $i$ is $C_{i}$.